\begin{thebibliography}{10}

\bibitem{li2024autoregressive}
Tianhong Li, Yonglong Tian, He~Li, Mingyang Deng, and Kaiming He.
\newblock Autoregressive image generation without vector quantization.
\newblock {\em Advances in Neural Information Processing Systems}, 37:56424--56445, 2024.

\bibitem{careil2023towards}
Marlene Careil, Matthew~J Muckley, Jakob Verbeek, and St{\'e}phane Lathuili{\`e}re.
\newblock Towards image compression with perfect realism at ultra-low bitrates.
\newblock In {\em The Twelfth International Conference on Learning Representations}, 2023.

\bibitem{poole2022dreamfusion}
Ben Poole, Ajay Jain, Jonathan~T Barron, and Ben Mildenhall.
\newblock Dreamfusion: Text-to-3d using 2d diffusion.
\newblock {\em arXiv preprint arXiv:2209.14988}, 2022.

\bibitem{sohl2015deep}
Jascha Sohl-Dickstein, Eric Weiss, Niru Maheswaranathan, and Surya Ganguli.
\newblock Deep unsupervised learning using nonequilibrium thermodynamics.
\newblock In {\em International conference on machine learning}, pages 2256--2265. PMLR, 2015.

\bibitem{ho2020denoising}
Jonathan Ho, Ajay Jain, and Pieter Abbeel.
\newblock Denoising diffusion probabilistic models.
\newblock {\em Advances in neural information processing systems}, 33:6840--6851, 2020.

\bibitem{song2020score}
Yang Song, Jascha Sohl-Dickstein, Diederik~P Kingma, Abhishek Kumar, Stefano Ermon, and Ben Poole.
\newblock Score-based generative modeling through stochastic differential equations.
\newblock {\em arXiv preprint arXiv:2011.13456}, 2020.

\bibitem{anderson1982reverse}
Brian~DO Anderson.
\newblock Reverse-time diffusion equation models.
\newblock {\em Stochastic Processes and their Applications}, 12(3):313--326, 1982.

\bibitem{vincent2011connection}
Pascal Vincent.
\newblock A connection between score matching and denoising autoencoders.
\newblock {\em Neural computation}, 23(7):1661--1674, 2011.

\bibitem{Song2019GenerativeMB}
Yang Song and Stefano Ermon.
\newblock Generative modeling by estimating gradients of the data distribution.
\newblock In {\em Neural Information Processing Systems}, 2019.

\bibitem{pan2025transfer}
Xichen Pan, Satya~Narayan Shukla, Aashu Singh, Zhuokai Zhao, Shlok~Kumar Mishra, Jialiang Wang, Zhiyang Xu, Jiuhai Chen, Kunpeng Li, Felix Juefei-Xu, et~al.
\newblock Transfer between modalities with metaqueries.
\newblock {\em arXiv preprint arXiv:2504.06256}, 2025.

\bibitem{deng2024autoregressive}
Haoge Deng, Ting Pan, Haiwen Diao, Zhengxiong Luo, Yufeng Cui, Huchuan Lu, Shiguang Shan, Yonggang Qi, and Xinlong Wang.
\newblock Autoregressive video generation without vector quantization.
\newblock {\em arXiv preprint arXiv:2412.14169}, 2024.

\bibitem{chen2024diffusion}
Boyuan Chen, Diego Mart{\'\i}~Mons{\'o}, Yilun Du, Max Simchowitz, Russ Tedrake, and Vincent Sitzmann.
\newblock Diffusion forcing: Next-token prediction meets full-sequence diffusion.
\newblock {\em Advances in Neural Information Processing Systems}, 37:24081--24125, 2024.

\bibitem{Yu2025FrequencyAI}
Hu~Yu, Hao Luo, Hangjie Yuan, Yu~Rong, and Feng Zhao.
\newblock Frequency autoregressive image generation with continuous tokens.
\newblock {\em ArXiv}, abs/2503.05305, 2025.

\bibitem{Han2025NextStep1TA}
NextStep Team~Chunrui Han, Guopeng Li, Jingwei Wu, Quan Sun, Yan Cai, Yuang Peng, Zheng Ge, Deyu Zhou, Haomiao Tang, Hongyu Zhou, Kenkun Liu, Ailin Huang, Bin Wang, Changxing Miao, Deshan Sun, En~Yu, Fukun Yin, Gang Yu, Hao Nie, Haoran Lv, Hanpeng Hu, Jia Wang, Jian Zhou, Jian‐Yuan Sun, Kaijun Tan, Kang An, Kangheng Lin, Liang Zhao, Mei Chen, Peng Xing, Rui Wang, Shiyu Liu, Shutao Xia, Tianhao You, Wei Ji, Xianfang Zeng, Xin pu~Han, Xuelin Zhang, Yana Wei, Yanming Xu, Yimin Jiang, Yingming Wang, Yu~Zhou, Yu~Han, Zi~Meng, Binxing Jiao, Daxin Jiang, Xiangyu~Tony Zhang, and Yibo Zhu.
\newblock Nextstep-1: Toward autoregressive image generation with continuous tokens at scale.
\newblock 2025.

\bibitem{Zhao2025DiSADS}
Qinyu Zhao, Jaskirat Singh, Ming Xu, Akshay Asthana, Stephen Gould, and Liang Zheng.
\newblock Disa: Diffusion step annealing in autoregressive image generation.
\newblock {\em ArXiv}, abs/2505.20297, 2025.

\bibitem{Fan2024FluidSA}
Lijie Fan, Tianhong Li, Siyang Qin, Yuanzhen Li, Chen Sun, Michael Rubinstein, Deqing Sun, Kaiming He, and Yonglong Tian.
\newblock Fluid: Scaling autoregressive text-to-image generative models with continuous tokens.
\newblock {\em ArXiv}, abs/2410.13863, 2024.

\bibitem{Tang2024HARTEV}
Haotian Tang, Yecheng Wu, Shang Yang, Enze Xie, Junsong Chen, Junyu Chen, Zhuoyang Zhang, Han Cai, Yao Lu, and Song Han.
\newblock Hart: Efficient visual generation with hybrid autoregressive transformer.
\newblock {\em ArXiv}, abs/2410.10812, 2024.

\bibitem{Meng2024AutoregressiveSS}
Lingwei Meng, Long Zhou, Shujie Liu, Sanyuan Chen, Bing Han, Shujie Hu, Yanqing Liu, Jinyu Li, Sheng Zhao, Xixin Wu, Helen Meng, and Furu Wei.
\newblock Autoregressive speech synthesis without vector quantization.
\newblock In {\em Annual Meeting of the Association for Computational Linguistics}, 2024.

\bibitem{yang2023lossy}
Ruihan Yang and Stephan Mandt.
\newblock Lossy image compression with conditional diffusion models.
\newblock {\em Advances in Neural Information Processing Systems}, 36:64971--64995, 2023.

\bibitem{Sargent2025FlowTT}
Kyle Sargent, Kyle Hsu, Justin Johnson, Fei-Fei Li, and Jiajun Wu.
\newblock Flow to the mode: Mode-seeking diffusion autoencoders for state-of-the-art image tokenization.
\newblock {\em ArXiv}, abs/2503.11056, 2025.

\bibitem{van2017neural}
Aaron Van Den~Oord, Oriol Vinyals, et~al.
\newblock Neural discrete representation learning.
\newblock {\em Advances in neural information processing systems}, 30, 2017.

\bibitem{wang2024taming}
Peihao Wang, Dejia Xu, Zhiwen Fan, Dilin Wang, Sreyas Mohan, Forrest Iandola, Rakesh Ranjan, Yilei Li, Qiang Liu, Zhangyang Wang, et~al.
\newblock Taming mode collapse in score distillation for text-to-3d generation.
\newblock In {\em Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 9037--9047, 2024.

\bibitem{Chung2024CFGMC}
Hyungjin Chung, Jeongsol Kim, Geon~Yeong Park, Hyelin Nam, and Jong~Chul Ye.
\newblock Cfg++: Manifold-constrained classifier free guidance for diffusion models.
\newblock {\em ArXiv}, abs/2406.08070, 2024.

\bibitem{mardani2023variational}
Morteza Mardani, Jiaming Song, Jan Kautz, and Arash Vahdat.
\newblock A variational perspective on solving inverse problems with diffusion models.
\newblock {\em arXiv preprint arXiv:2305.04391}, 2023.

\bibitem{alkan2023variational}
Cagan Alkan, Julio Oscanoa, Daniel Abraham, Mengze Gao, Aizada Nurdinova, Kawin Setsompop, John~M Pauly, Morteza Mardani, and Shreyas Vasanawala.
\newblock Variational diffusion models for blind mri inverse problems.
\newblock In {\em NeurIPS 2023 workshop on deep learning and inverse problems}, 2023.

\bibitem{Pandey2025VariationalCF}
Kushagra Pandey, Farrin~Marouf Sofian, Felix Draxler, Theofanis Karaletsos, and Stephan Mandt.
\newblock Variational control for guidance in diffusion models.
\newblock {\em ArXiv}, abs/2502.03686, 2025.

\end{thebibliography}

\documentclass{article}

 \usepackage[dblblindworkshop, final]{neurips_2025}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{xcolor}         
\usepackage{hyperref}
\usepackage{url}
\usepackage{hyperref}
\usepackage{url}
\usepackage{bbm}
\usepackage{amsthm}
\newtheorem{theorem}{Theorem}
\newtheorem{corollary}{Corollary}
\usepackage{booktabs}
\usepackage{multirow}
\usepackage{graphics}
\usepackage{graphicx}
\usepackage{booktabs}
\usepackage{multirow}
\usepackage{caption}
\usepackage{etoolbox}
\usepackage{makecell}
\usepackage{listings}

\usepackage{hyperref}
\usepackage{url}
\usepackage{amsthm}
\usepackage{graphicx}
\usepackage{pifont}
\usepackage{bbm}
\usepackage{bm}
\usepackage{caption}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathtools}

\theoremstyle{definition}

\usepackage{multirow}
\usepackage{adjustbox}
\usepackage[linesnumbered,ruled]{algorithm2e}
\usepackage{booktabs}
\usepackage{amssymb}
\usepackage{dirtytalk}
\usepackage{xurl}
\usepackage{thmtools}
\usepackage{thm-restate}
\usepackage{arydshln}
\usepackage{tablefootnote}
\usepackage{dirtytalk}
\usepackage{listings}
\usepackage{xcolor}

\title{Optimizing Input of Denoising Score Matching is Biased Towards Higher Score Norm}
\workshoptitle{Frontiers in Probabilistic Inference: Sampling Meets Learning}
\author{%
  Tongda Xu \\
  AIR \& CST, Tsinghua University\\
  \texttt{x.tongda@nyu.edu} \\
}

\begin{document}

\maketitle

\begin{abstract}
Many recent works utilize denoising score matching to optimize the conditional input of diffusion models. In this workshop paper, we demonstrate that such optimization breaks the equivalence between denoising score matching and exact score matching. Furthermore, we show that this bias leads to higher score norm. Additionally, we observe a similar bias when optimizing the data distribution using a pre-trained diffusion model. Finally, we discuss the wide range of works across different domains that are affected by this bias, including MAR for auto-regressive generation \citep{li2024autoregressive}, PerCo \citep{careil2023towards} for image compression, and DreamFusion \citep{poole2022dreamfusion} for text to 3D generation.
\end{abstract}

\section{Introduction}
The diffusion model is a type of generative model that consists of a Gaussian Markov chain in continuous space \citep{sohl2015deep,ho2020denoising,song2020score}. We denote the source image as $x_0$, and the forward diffusion process involves perturbing the image with Gaussian noise, i.e., $q(x_t) = x + \mathcal{N}(0, \sigma^2_t)$. To sample from the backward diffusion model, it is necessary to know the score function $\nabla_{x_t} \log q(x_t)$ in order to construct the reverse stochastic differential equation \citep{anderson1982reverse}.

The learning of diffusion models relies on the estimation of the score function $\nabla_{x_t}\log q(x_t)$. The most straightforward approach is explicit score matching (ESM), which employs a neural network $s_{\theta}(x_t, t)$ parameterized by $\theta$ to approximate the score:
\begin{gather}
   \mathcal{L}_{ESM}(\theta) = \mathbb{E}_{q(x_t|x)p(x)}[||s_{\theta}(x_t,t) - \nabla_{x_t} \log q(x_t)||^2].
\end{gather}
However, $\mathcal{L}_{ESM}$ is intractable because $\nabla{x_t}\log q(x_t)$ is itself intractable. \citep{vincent2011connection,Song2019GenerativeMB} proposed using denoising score matching (DSM) instead, as $\nabla_{x_t}\log q(x_t|x)$ is simply the score of a Gaussian distribution. Furthermore, \citep{vincent2011connection} proved the equivalence between the $\mathcal{L}_{ESM}$ objective and the $\mathcal{L}_{DSM}$ objective when optimizing $\theta$:
\begin{gather}
    \mathcal{L}_{DSM}(\theta) = \mathbb{E}_{q(x_t|x)p(x)}[||s_{\theta}(x_t,t) - \nabla_{x_t} \log q(x_t|x)||^2].
\end{gather}

\begin{theorem} \citep{vincent2011connection} When optimizing $\theta$, the denoising score matching is equivalent to explicit score matching, \textit{i.e.}, 
\begin{gather}
    \mathcal{L}_{DSM}(\theta) \smallsmile \mathcal{L}_{ESM}(\theta).
\end{gather}
\end{theorem}
Many recent works have begun to use $\mathcal{L}_{DSM}$ to optimize targets other than $\theta$. For example, MAR \citep{li2024autoregressive} adopts $\mathcal{L}_{DSM}$ to optimize $c$, the conditional input of the diffusion model to train an auto-regressive image generation model:
\begin{gather}
    \theta^*,c^* \leftarrow \arg\max \mathcal{L}_{DSM}(\theta,c) = \mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||s_{\theta}(x_t,t,c) - \nabla_{x_t}\log q(x_t|x,c)||^2].
\end{gather}
Another example is SDS \citep{poole2022dreamfusion} for text-to-3D generation, which utilizes a pre-trained diffusion model $s_{\theta}(.,.,.)$ and $\mathcal{L}_{DSM}$ to optimize the input distribution $p(x)$:
\begin{gather}
    p(x)^* \leftarrow \arg\max \mathcal{L}_{DSM}(p(x)) = \mathbb{E}_{q(x_t|x)p(x)}[\frac{1}{2}||s_{\theta}(x_t,t) - \nabla_{x_t}\log q(x_t|x)||^2].
\end{gather}
In this workshop paper, we show that the above usage of $\mathcal{L}_{DSM}$ is biased, as it breaks the equivalence between $\mathcal{L}_{DSM}$ and $\mathcal{L}_{ESM}$. Moreover, we demonstrate that this bias drives the optimization towards higher score norm. Finally, we discuss the broad range of works that are potentially affected by this bias.

\section{Optimizing Condition with Diffusion Loss is Biased Towards Higher Score Norm}
According to \citep{vincent2011connection}, $\mathcal{L}_{DSM}$ and $\mathcal{L}_{ESM}$ have the following relationship:
\begin{align}    
    \mathcal{L}_{DSM}(\theta,c) =& \mathcal{L}_{ESM}(\theta,c) - \underbrace{\mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2]}_{C_2} \notag \\&+ \underbrace{\mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|x,c)||^2]}_{C_3}.
\end{align}
When optimizing $\theta$, the $C_2$ and $C_3$ terms cancel out, as they are not related to $\theta$. However, although $\theta$ is not related to these terms, the condition $c$ is involved in both the $C_2$ and $C_3$ terms. Since $x_t$-$x$-$c$ forms a Markov chain, the $C_3$ term has no gradient with respect to $c$.

Therefore, when optimizing $c$ using $\mathcal{L}_{DSM}$, there is an additional term compared to $\mathcal{L}_{ESM}$. Since the equivalence between $\mathcal{L}_{DSM}$ and $\mathcal{L}_{ESM}$ is broken, $\mathcal{L}_{DSM}$ becomes biased.

\begin{theorem} When optimizing both $\theta$ and $c$, the denoising score matching is equivalent to explicit score matching with a bias term, \textit{i.e.},
\begin{gather}
    \mathcal{L}_{DSM}(\theta,c) \smallsmile \mathcal{L}_{ESM}(\theta,c) - \underbrace{\mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2]}_{C_2}.
\end{gather}
\end{theorem}
We have shown that for the condition $c$, minimizing $\mathcal{L}_{DSM}$ is equivalent to simultaneously minimizing $\mathcal{L}_{ESM}$ and maximizing $C_2 = \mathbb{E}{q(x_t|x) p(x|c)}\left[\frac{1}{2}||\nabla{x_t}\log q(x_t|c)||^2\right]$. And therefore, optimizing $\mathcal{L}_{DSM}$ is biased towards higher score norm.

\section{Optimizing Data Distribution with Diffusion Loss is Also Biased Towards Higher Score Norm}
A similar result can be trivially shown when using the score matching diffusion loss with a pre-trained diffusion model to optimize the input data distribution $p(x)$.
\begin{corollary}
When optimizing input distribution $p(x)$ with pre-trained diffusion model, the denosiing score matching is equivalent to explicit score matching with a bias term, \textit{i.e.,}
\begin{align}
    \mathcal{L}_{DSM}(p(x)) \smallsmile \mathcal{L}_{ESM}(p(x)) - \mathbb{E}_{q(x_t|x)p(x)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t)||^2].
\end{align}
\end{corollary}
\section{Discussion: Works That are Effected by This Bias}
Optimizing the condition or input distribution using the diffusion loss is a common technique in the application of diffusion models across various domains. Below, we highlight four areas with recent works that are affected by this bias; this list is by no means exhaustive.

\textbf{Auto-regressive Generation}: Various works in tokenization-free auto-regressive generation are affected by this bias. MAR \citep{li2024autoregressive} adopts the diffusion loss to optimize $c$, which is the output of the auto-regressive model. MetaQuery \citep{pan2025transfer} employs diffusion loss to optimize $c$, which is the output of the connector in MLLM models. Similarly, \citep{deng2024autoregressive,chen2024diffusion,Yu2025FrequencyAI,Han2025NextStep1TA,Zhao2025DiSADS,Fan2024FluidSA,Tang2024HARTEV,Meng2024AutoregressiveSS} are also effected by this bias.

\textbf{Image Compression}: CDC \citep{yang2023lossy} jointly minimizes the bitrate of $c$ alongside the diffusion loss. PerCo \citep{careil2023towards} optimizes $c$ with the diffusion loss, where the resulting gradient is used to refine a VQ-VAE. FlowMo \citep{Sargent2025FlowTT} optimizes $c$ with diffusion loss, whose gradient is used to jointly train a VQ-VAE \citep{van2017neural}.

\textbf{Text-to-3D Generation}: DreamFusion \citep{poole2022dreamfusion} utilizes a pre-trained 2D diffusion model to optimize the input distribution $p(x)$ using the diffusion loss. The gradient with respect to $p(x)$ is then used to optimize a 3D representation model that is employed to render $p(x)$. Similarly, \citep{wang2024taming,Chung2024CFGMC} are also effected by this bias.

\textbf{Diffusion-based Inverse Problem Solving}: RED-Diff \citep{mardani2023variational} adopts the loss function used in DreamFusion to optimize $p(x)$, thereby solving inverse problems for operators such as super-resolution and de-blurring. Similarly, \citep{alkan2023variational,Pandey2025VariationalCF} are also effected by this bias.
\medskip
{
\small
\bibliographystyle{unsrt}
\bibliography{main}
}


\appendix

\section{Proof of Main Results}

\textbf{Theorem 2. }\textit{When optimizing both $\theta$ and $c$, the denoising score matching is equivalent to explicit score matching with a bias term, \textit{i.e.},
\begin{gather}
    \mathcal{L}_{DSM}(\theta,c) \smallsmile \mathcal{L}_{ESM}(\theta,c) - \underbrace{\mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2]}_{C_2}.
\end{gather}
}
\begin{proof}
    \begin{align}
    \mathcal{L}_{ESM}(\theta,c)=&\mathbb{E}_{q(x_t|c)}[\frac{1}{2}||s_{\theta}(x_t,t,c) - \nabla_{x_t}\log q(x_t|c)||^2] \notag \\
    =& \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||s_{\theta}(x_t,t,c)||^2] - \underbrace{\mathbb{E}_{q(x_t|c)}[s_{\theta}(x_t,t,c)^T\nabla_{x_t}\log q(x_t|c)]}_{S_1} \notag \\ &+ \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2]. \label{eq:esm}
    \end{align}
According to \citep{vincent2011connection}, the $S_1$ term can be transformed by applying log derivative twice:
    \begin{align}
        S_1 =& \mathbb{E}_{q(x_t|c)}[s_{\theta}(x_t,t,c)^T\nabla_{x_t}\log q(x_t|c)] \notag \\
        =&\int_{x_t}q(x_t|c) s_{\theta}(x_t,t,c)^T\nabla_{x_t}\log q(x_t|c) dx_t \notag \\
        =&\int_{x_t}q(x_t|c) s_{\theta}(x_t,t,c)^T\frac{\nabla_{x_t} q(x_t|c)}{q(x_t|c)} dx_t \notag \\
        =& \int_{x_t} s_{\theta}(x_t,t,c)^T\nabla_{x_t} q(x_t|c) dx_t \notag \\
        =& \int_{x_t} s_{\theta}(x_t,t,c)^T(\nabla_{x_t}\int_{x}p(x|c)q(x_t|x,c)dx) dx_t \notag \\
        =& \int_{x_t} s_{\theta}(x_t,t,c)^T(\int_{x}p(x|c)\nabla_{x_t}q(x_t|x,c)dx) dx_t \notag \\
        =& \int_{x_t} s_{\theta}(x_t,t,c)^T(\int_{x}p(x|c)q(x_t|x,c)\nabla_{x_t}\log q(x_t|x,c)dx) dx_t \notag \\
        =& \int_{x_t}\int_{x}p(x|c)q(x_t|x,c) s_{\theta}(x_t,t,c)^T(\nabla_{x_t}\log q(x_t|x,c)dx) dx_t \notag \\
        =& \mathbb{E}_{q(x_t|x)p(x|c)} [s_{\theta}(x_t,t,c)^T\nabla_{x_t}\log q(x_t|x,c)] \label{eq:s1}.
    \end{align}
Substituting Eq.~\ref{eq:s1} into Eq.~\ref{eq:esm}, we have
\begin{align}
    \mathcal{L}_{ESM}(\theta,c)=& \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||s_{\theta}(x_t,t,c)||^2] - \mathbb{E}_{q(xt,x|c)} [s_{\theta}(x_t,t,c)^T\nabla_{x_t}\log q(x_t|x,c)] \\ \notag &+ \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2].
\end{align}
On the other hand, we have
\begin{align}
    \mathcal{L}_{DSM}(\theta,c) =& \mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||s_{\theta}(x_t,t,c) - \nabla_{x_t}\log q(x_t|x,c)||^2] \notag \\
    =& \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||s_{\theta}(x_t,t,c)||^2] - \mathbb{E}_{q(xt,x|c)} [s_{\theta}(x_t,t,c)^T\nabla_{x_t}\log q(x_t|x,c)] \notag \\
    &+ \underbrace{\mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|x,c)||^2]}_{C_3} \notag \\
    =& \mathcal{L}_{ESM}(\theta,c) - \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2] + C_3.
\end{align}
For the $C_3$ term, we notice that $x_t-x-c$ forms a Markovian chain, and therefore 
\begin{align}
    C_3 = \mathbb{E}_{q(x_t|x)p(x|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|x)||^2].
\end{align}
As $C_3$ is not related to $\theta$ and $c$, we have
\begin{gather}
    \mathcal{L}_{ESM}(\theta,c) - \mathbb{E}_{q(x_t|c)}[\frac{1}{2}||\nabla_{x_t}\log q(x_t|c)||^2] \smallsmile \mathcal{L}_{DSM}(\theta,c).
\end{gather}
\end{proof}


\end{document}